0.50 g of an organic compound containing carbon, hydrogen and oxygen gives 0.6875 g of carbon dioxide and 0.5625 g of water on combustion. Find the empirical formula of the compound.
(Relative atomic mass: H:1; C: 12; O: 16)
1
Expert's answer
2020-09-21T06:34:20-0400

CxHyOz + O2 → CO2 + H2O

m(CxHyOz) = 0.5 g

m(CO2) = 0.6875 g

m(H2O) = 0.5626 g

M(CO2) = 44 g/mol

M(H2O) = 18 g/mol

n = m/M

n(CO2) = 0.6875/44 = 0.0156 mol

Since one mole of CO2 is made up of one mole of C and two moles of O, if we have 0.0156 moles of CO2 in our sample, then we know we have 0.0156 moles of C in the sample.

m(C) = 0.0156*12 = 0.1872 g

n(H2O) = 0.5626/18 = 0.0312 mol

Since one mole of H2O is made up of one mole of oxygen and two moles of hydrogen, if we have 0.0312 moles of H2O, then we have 2*(0.0312) = 0.0624 moles of hydrogen. Since hydrogen is about 1 gram/mole, we must have 0.0624 grams of hydrogen in our original sample.

0.1872 g of C + 0.0624 g of H = 0.2496 g

m(O) = 0.5 – 0.2496 = 0.2504 g

n(O) = 0.2504/16 = 0.0156 mol

Proportion: C:H:O = 0.0156:0.0624:0.0156=1:4:1

The empirical formula for an organic compound would be CH4O.

Answer: CH4O.